OpenStudy (anonymous):

Factor out the common factor of 5a. 20a^9/4-25a^8/5

OpenStudy (anonymous):

go to http://www.freemathhelp.com/

OpenStudy (anonymous):

\[5a(4a^{5/4}-5a^{3/5})\]

OpenStudy (anonymous):

I got it wrong?

OpenStudy (anonymous):

What did you get?

OpenStudy (anonymous):

Nevermind....I messed up, you were right.....Sorry

OpenStudy (anonymous):

\[5a(4a^{5/4}-5a^{3/5})=5a \times 4a^{4/5} - 5a \times 5a^{3/5}\]\[=5a^{4/4} \times 4a^{5/4}-5a^{5/5} \times 5a^{3/5}\]=\[20a^{\frac{4+5}{4}}-25^{\frac{5+3}{5}}=20^{9/4}-25^{8/5}\]